OpenStudy (goformit100):

A grandfather is ten times older than his granddaughter. He is also 54 years older than her. Find their present ages.

OpenStudy (goformit100):

@Justin9529

OpenStudy (justin9529):

Grandfather’s age = Granddaughter’s age + 54 years 10x = x + 54 Transposing x to L.H.S, we obtain 10x − x = 54 9x = 54 x = 6 Granddaughter’s age = x years = 6 years Grandfather’s age = 10x years = (10 × 6) years = 60 years

OpenStudy (nincompoop):

g = 10d 10d = d+54 9d = 54 d = 6 g = 10d g = 10(6) g = 60 where d is daughter and g is grandfather

OpenStudy (goformit100):

thankyousir